OpenStudy (anonymous):

a company uses two vans to transport workers from a free parking lot to the workplace between 7:00 and 9:00 a.m. one van has 14 more seats than the other. the smalker van makes two trips every morning while the larger one makes only one the two vans can transport 65 people maximum.

OpenStudy (anonymous):

how many seats does a larger van have @KamiBug

OpenStudy (shinalcantara):

Let 'x' be the capacity of the smaller van. 'x+14' be the capacity of the larger van ------------------------- Since it is stated that the smaller van makes two trips every morning, then we can say that the total passengers the van was able to transport is twice its capacity or 2x. ------------------------- The two vans can transport 65 people then, 2x + x + 14 = 65 transpose 14 to the right 2x + x = 65 - 14 3x = 51 x = 17 ----------------------- Since the larger van's capacity is x + 14 then, x + 14 = 17 + 14 = 31 ---------------------- Therefore there are 31 seats in the larger van